wiener:

Please help me with some math questions:

wiener:

wiener:

AP:

6.) Since AC and AB are congruent, it means that <ACB = <ABC That means m<B also equals 69° ~~> 69 + 69 + m<A = 180°

wiener:

ok thanks

AP:

So m<A equals 42°

wiener:

yes

AP:

7.) m<D = m<F So m<D + m<F + 77° = 180° 180 - 77 = 103 103 / 2 = 51.5° That means m<D = 51.5° and m<F=51.5°

wiener:

ok thanks

wiener:

makes sense

AP:

8.) GJ = HJ 3x + 8 = 35 and solve for x

SUPAflyGUY:

thank you

SUPAflyGUY:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @wiener Please help me with some math questions: \(\color{#0cbb34}{\text{End of Quote}}\) also posted by me

AP:

3x + 8 = 35 -8 -8 3x = 27 /3 /3 x = 9

SUPAflyGUY:

ok

AP:

To find the measure of GH, substitute 9 in for x in the equation 6x - 5 @SUPAflyGUY

SUPAflyGUY:

ok uno momento

SUPAflyGUY:



AP:

You got it GH = 49°

AP:

Not degrees lol typo

SUPAflyGUY:

lol ok

AP:

9) LK = LM So 26 = 2y + 8 Solve for y @SUPAflyGUY

SUPAflyGUY:

ok @AP

SUPAflyGUY:

y=9 @AP

AP:

Good. Now plug in 9 for y in the equation 3y - 7 to solve for KM

SUPAflyGUY:

=9

AP:

No, like \[3(9) - 7 = ?\]

SUPAflyGUY:

20

AP:

You got it (:

SUPAflyGUY:

what about 10? @AP

AP:

You didnt post 10.)

SUPAflyGUY:

AP:

C = A Means both angles are 71° So m<ABD = 180 - 90 - 71 Solve @SUPAflyGUY

SUPAflyGUY:

19?

SUPAflyGUY:

@AP

AP:

Correct

SUPAflyGUY:

ok thats the answer?

AP:

That's it

SUPAflyGUY:

ok thanks a lot!

AP:

You're welcome (: